Circular No. 7961

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 Q4 (BREWINGTON) F. Artigue, H. Cucurullo, and G. Trancredi, Observatorio Astronómico

Los Molinos, Montevideo, report the recovery of comet P/1992 Q1 (= 1992p = 1992 XIV), with a diffuse coma of diameter 20′′ and central condensation, on CCD images taken with a 0.46-m telescope in the course of the ‘BUSCA’ project. Further astrometry and orbital elements (from observations 1992 Aug. 28–2002 Aug. 28) appear on MPEC 2002-Q41; the correction to the prediction on MPC 40670 is ΔT = +0.52 day.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 26.97571  20^{h}07^{m}01.^{s}59  −36^{o}10′07′′.1
27.98148  20 06 06.51  −36 02 34.2  16.8

SUPERNOVA 2002er IN UGC 10743 S. J. Smartt, Institute of Astronomy, University of Cambridge; F.

Patat, European Southern Observatory; and P. Meikle, Imperial College, London, on behalf of the European Research and Training Network on the Physics of Type-Ia Supernovae, report that a 1800-s spectrum (range 360– 720 nm; resolution 0.46 nm) of SN 2002er (cf. IAUC 7959), obtained at the Isaac Newton Telescope (+ IDS spectrograph) on La Palma by S. Araujo on Aug. 26.93 UT, shows it to be a type-Ia supernova. A comparison with the spectra of SN 1994D (Patat et al. 1996, MNRAS 278, 111), and in particular the relative strengths of the features of Fe II/III and Si II, suggests an early epoch of roughly 10 days prior to peak brightness. The spectrum is significantly redder than that of SN 1994D, indicating host-galaxy extinction. The interstellar Na I D lines from the host galaxy at +2508 km/s are seen as an unresolved blend in the spectrum and have equivalent widths of 0.11 ± 0.01 nm for the Milky Way and 0.12 ± 0.01 nm for UGC 10743. The Galactic extinction of Schlegel et al. (1998, Ap.J. 500, 525) suggests a reddening of E(B − V) = 0.16 towards UGC 10743. Since the Na I lines of the host-galaxy absoprtion are similar to the Milky Way absorption, the likely total reddening is of the order E(B − V) ∼ 0.3.

SUPERNOVA 2002eo IN NGC 710 M. Armstrong, Rolvenden, Kent, England, reports the following mag-

nitudes for SN 2002eo (cf. IAUC 7958) from unfiltered CCD frames taken in the course of his supernova search: 2000 Sept. 25 UT, [19.5; 2002 Aug. 14.077, [19.5; 18.023, 18.5: (faint trace; poor image); 23.089, 18.2:.

2002 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green